OpenStudy (anonymous):

the angles of from smallest to largest if AB = 21, BC = 14, and AC = 8.

OpenStudy (anonymous):

Angles of triangle ABCD from smallest to largest, help please

OpenStudy (triciaal):

the largest angle is opposite the longest side the smallest angle is opposite the smallest side

OpenStudy (triciaal):

how can you have a triangle ABCD?

RhondaSommer (rhondasommer):

I can help

RhondaSommer (rhondasommer):

do you know how to do sin cos tan?

RhondaSommer (rhondasommer):



RhondaSommer (rhondasommer):

tan=8/14 = b tan=14/8 = a 90-a-b = c

OpenStudy (triciaal):

@RhondaSommer you do not have a right angle triangle the sum of the angles in a triangle is 180 degrees

RhondaSommer (rhondasommer):

whoops your right lol...meant to put 180 :P

RhondaSommer (rhondasommer):

you're*

RhondaSommer (rhondasommer):

so 180-a-b = c

OpenStudy (anonymous):

I meant abc triangle sorry

OpenStudy (triciaal):

do you need to calculate the angles?

OpenStudy (triciaal):

@RhondaSommer and @Kalin234 when only given the sides of the triangle use the cosine rule

RhondaSommer (rhondasommer):

she would if she has to get the angles from largest to smallest or smallest to largest

RhondaSommer (rhondasommer):

you dont have to use cos though

OpenStudy (triciaal):

after the largest angle then you can use the sin rule to find the others

OpenStudy (anonymous):

I have to organize the angles from largest and smalles

OpenStudy (anonymous):

Smallest to largest

OpenStudy (triciaal):

review the thread

OpenStudy (anonymous):

Bc is the smallest

OpenStudy (anonymous):

Or ac

RhondaSommer (rhondasommer):

that is the smallest side not angle

OpenStudy (triciaal):